# Lite-Mind: Towards Efficient and Robust Brain Representation Learning

### v2

@@ 3.2 DFT Backbone
$\begin{split}X^{\prime}&=({\sigma(Re(\hat{X}^{T})\mathcal{W}_{r}-Im(\hat{X}^{T})\mathcal{W}_{i}+\mathcal{B}_{r})}\\ &+j\sigma(Re(\hat{X}^{T})\mathcal{W}_{i}+Im(\hat{X}^{T})\mathcal{W}_{r}+\mathcal{B}_{i}))^{T}\end{split}$ (5)
Where $Re(\cdot)$ denotes the real part of $\hat{X}^{T}$, $Im(\cdot)$ denotes the imaginary part, $\mathcal{W}=\mathcal{W}_{r}+j\mathcal{W}_{i}$ and $\mathcal{B}=\mathcal{B}_{r}+j\mathcal{B}_{i}$. Due to the more significant features of filtered and compressed tokens in the frequency domain, we replace the final FC layer of MindEye’s MLP Backbone with FreMLP. For more theoretical proof and implementation of FreMLP, please refer to the Appendix D.2. Finally, we follow IDFT $t^{\prime}\leftarrow F^{-1}(X^{\prime})$ to get final tokens $t^{\prime}$, and use voxel embedding $f$ to represent them.
3.3 Retrieval Pipeline
3.3.1 Optimization Objective
The optimization objective for the retrieval task is defined as follows:
@@ 3.3.1 Optimization Objective
3.3.2 Diffusion Projector
When searching on a large-scale dataset like LAION-5B, image-to-image performs better than fMRI-to-image and can find more similar images. Therefore, we used a diffusion projector to translate the disjointed voxel CLS embeddings of the DFT Backbone output, using a DALLE · 2[31] similar diffusion model to obtain the image CLS embeddings and retrieve it on LAION-5B: $\mathcal{V^{\prime}}=Diffusion(f)$. The Diffusion Projector consists of a transformer for image CLS embedding forecasting.
3.3.3 Contrastive Learning
@@
$L_{contr}=-\frac{1}{|B|}\sum_{s=1}^{|B|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|B|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$ (7)
@@
$L_{mse}=\frac{1}{|B|}\sum_{s=1}^{|B|}{||\mathcal{V}_{s}-\mathcal{V^{\prime}}_{s}||}^{2}_{2}$ (8)
@@ 3.3.4 Downstream tasks
Figure 3: Partial retrieval results of Lite-Mind on all 982 test images for Subject 1. With 12.5M DFT Backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates. The number below each image represents the similarity score. See more cases include failure retrieval in Appendix C.4.
4 Experiments
4.1 Dataset
Natural Scenes Dataset (NSD) is an extensive 7T fMRI dataset gathered from 8 subjects viewing images from MSCOCO-2017 dataset [1], which contains images of complex natural scenes. Participants viewed three repetitions of 10,000 images with a 7-Tesla fMRI scanner over 30–40 sessions. More details can be found on the NSD official website [^1] https://naturalscenesdataset.org. Our experiments focused on Subj01, Subj02, Subj05, and Subj07, who finished all viewing trials. Each subject’s training set comprises 8859 image stimuli and 24980 fMRI trials (with the possibility of 1-3 repetitions per image). The test set contains 982 image stimuli and 2770 fMRI trials. Responses for images with multiple fMRI trials are averaged across these trials. By applying the nsdgeneral ROI mask with a 1.8 mm resolution, we obtained ROIs for the four subjects, comprising 15724, 14278, 13039, and 12682 voxels respectively. These regions span visual areas ranging from the early visual cortex to higher visual areas. Our experimental setup is consistent with the NSD image reconstruction and retrieval articles [38, 27, 25, 35, 21].
Generic Object Decoding (GOD) Dataset [^2] http://brainliner.jp/data/brainliner/Generic_Object_Decoding was created by Horikawa and Kamitani [13], consisting of fMRI recordings of five healthy subjects who were presented with images from ImageNet [6]. The GOD Dataset includes 1250 distinct images selected from 200 ImageNet categories. Among these, 1200 training images are drawn from 150 categories, and 50 test images are from the remaining 50 categories. The training and test image stimuli were presented to the subjects once and 35 times respectively, resulting in 1200 and 1750 fMRI instances. We used preprocessed ROIs, encompassing voxels from the early visual cortex to higher visual areas. For each test image, the fMRI responses from different trials were averaged.
4.2 Implementation details
@@
5 Results
5.1 fMRI/image retrieval
Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings on the test set. If a paired image embedding is retrieved, the retrieval is considered correct. fMRI retrieval is the opposite process mentioned above. Note that there are many semantically and visually similar images on the NSD dataset test set, which are considered to be similar in CLIP space. Whether the model can correctly retrieve as MindEye tests the fine-grained alignment ability to the image. For test retrieval, we adhered to the identical methodology as Lin et al. [21] and MindEye [35] to compute the retrieval metrics presented in Table 0(a). For each test sample, we randomly selected 299 images from the remaining 981 images in the test set and calculated the cosine similarity between the voxel embeddings and 300 images. The retrieval accuracy refers to the proportion of successful retrieval of corresponding images in the 982 voxel embeddings of the test set. We adjusted the random number seed of 30 randomly selected images to average the accuracy of all samples. The experimental results are shown in Table 0(a). We also conducted retrieval experiments on the remaining three subjects to demonstrate the universality of DFT Backbone. The detailed results and discussions are in Appendix B.2.
@@ 5 Results
Lin et al. [21] deep models 2.34M 11.0% 49.0%
Ozcelik… [27] 257 separate linear regression models 3B 29.9% 21.4%
MindEye [35] MLP Backbone 940M 89.6% 82.2%
MindEye [35] MLP Backbone+Projector 996M 88.8% 84.9%
MindEye [35] MLP Backbone+Prior 1.2B 93.4% 90.1%
@@
Lin et al. [21] - - - - 78.2% - - -
Takagi… [38] - - 83.0% 83.0% 76.0% 77.0% - -
Gu et al. [10] .150 .325 - - - - .862 .465
Ozcelik… [27] .254 .356 94.2% 96.2% 87.2% 91.5% .775 .423
BrainCLIP [23] - - - - 86.7% 94.8% - -
MindEye [35] .309 .323 94.7% 97.8% 93.8% 94.1 .645 .367
@@
Table 1: Main retrieval results on the NSD and LAION-5B dataset.
5.2 LAION-5B retrieval
@@
Although the retrieval scale on NSD is large enough, we can still expand it to larger datasets, such as LAION-5B [34]. The final layer CLIP ViT-L/14 CLS embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client [2]. For LAION-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS embeddings of CLIP. For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS embeddings, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP). Unlike MindEye, in the LAION-5B retrieval process, we used the Diffusion Projector to conduct image-to-image retrieval.
@@
To compare the performance of large-scale dataset retrieval, we use metrics of image reconstruction for evaluation with previous image reconstruction methods. The experimental results are shown in Table 0(b). The specific evaluation metrics in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstruction images; SSIM is structural similarity index metric [40]; EfficientNetB1(”Eff”) [39] and SwAV-ResNet50(”SwAV”) [4] refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
Experiments show that our DFT Backbone performs better than MindEye in high-level metrics, as shown in Table 0(b), although our backbone only has 8M parameters(MindEye-CLS backbone still has 125M. Note that both MindEye and Lite-Mind perform LAION-5B retrieval based on CLS alignment, resulting in a smaller number of parameters compared to full models). The results indicate that an efficient DFT backbone can also assist downstream retrieval tasks. Partial retrieval visualization results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval and visualization results of each Subject are shown in Appendix B.2.
5.3 GOD zero-shot classification
Since the collective fMRI data are much fewer and shorter than those of the NSD dataset usually, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify the Lite-Mind’s generalization ability. The voxel lengths collected by Subjects on the GOD dataset range from 4133 to 4643, approximately 30% of NSD, and there are only 1200 fMRI-image pairs on the training set (4.8% of NSD).
We conduct zero-shot classification by using fMRI-to-image retrieval to find the specific image and using simple prompt text templates ”An image of [class]” to obtain the category of retrieved images(CLIP has a classification accuracy of 76.2% on ImageNet without pre-training on it, so it can be seen as zero-shot classification). We train our Lite-Mind-15.5M to align voxels to images on the GOD dataset. We also replicated the MLP backbone and projector of MindEye, changing the residual block of $4096\times 4096$ to $1024\times 1024$ to accommodate the reduction of voxel length(We empirically found that the original $4096\times 4096$ performed even worse), and MindEye still has 210M parameters. As shown in Table 2, the experimental results indicate that our proposed Lite-Mind still guarantees robustness and establishes a new state-of-the-art for zero-shot classification on the GOD dataset without complex text prompt templates or auxiliary text representation training. On the contrary, due to the sudden decrease in dataset size and voxel length, the performance of MindEye significantly decreases, indicating excessive reliance on training data size for voxel value-wise mappings with large MLP Backbone. More retrieval and classification results are in Appendix B.3.
@@ 5.3 GOD zero-shot classification
Figure 4: Partial retrieval results of Lite-Mind from LAION-5B for Subject 1. The left column marked by a red box in every two columns represents the original image seen by the Subject, and the right column represents the image retrieved from LAION-5B.
@@
CADA-VAE [33] V&T - 6.31 35.70 6.45 40.12 17.74 54.34 12.17 36.64 7.45 35.04
MVAE [41] V&T - 5.77 31.51 5.40 38.46 17.11 52.46 14.02 40.90 7.89 34.63
MMVAE [37] V&T - 6.63 38.74 6.60 41.03 22.11 56.28 14.54 42.45 8.53 38.14
MoPoE-VAE [24] V&T - 8.54 44.05 8.34 48.11 22.68 61.82 14.57 58.51 10.45 46.40
BraVL [7] V&T - 9.11 46.80 8.91 48.86 24.00 62.06 15.08 60.00 12.86 47.94
BrainClip-VAE [23] V&T Text 8.00 42.00 24.00 52.00 20.00 58.00 20.00 58.00 20.00 46.00
BrainClip-VAE [23] V&T CoOp 14.00 54.00 20.00 60.00 21.33 64.67 16.67 66.67 18.00 52.00
MindEye-211M [35] V Text 6.00 40.00 20.00 60.00 28.00 70.00 14.00 54.00 10.00 64.00
@@
Table 2: Zero-shot visual stimulus classification on the GOD dataset. The test set contains 50 categories that have no overlapping with the training set (top-1 chance=2.0%). The results for CADA-VAE, MVAE, MMVAE, MoPoE-VAE, and BraVL are taken from [7]. MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B.3 for retrieval performance on the GOD dataset).
5.4 Ablations and visualization
@@
Architectural Analysis. We trained multiple DFT Backbone models with different depths to evaluate the impact of the number of layers of fMRI Spectrum Compression on retrieval efficiency(Table 3). As the Filter Blocks layers gradually deepen, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual. We interestingly found that the accuracy of brain retrieval tends to converge faster than the image retrieval and only 6 layers of Filter Blocks are needed to approach the highest value. We have also conducted more ablation experiments on hyper-parameters to demonstrate the robustness of Lite-Mind. Refer to Appendix B.4 for it.
@@
Table 3: Ablation experiments on fMRI Spectrum Compression revealed a positive correlation between the depth of the Filter Blocks and enhanced forward retrieval accuracy.
Module Analysis. We conducted ablation experiments on two main modules of DFT Backbone in Table 4. We completely removed the Filter blocks and aligned tokens directly to the image embeddings with fMRI token decoupling, achieving a forward retrieval accuracy of 65.8%, indicating that the importance of frequency domain filtering for denoising and compression. Then we replaced FreMLP with the FC layer in the real domain, where tokens are flatten into a tensor for alignment, similar to the last layer of the MLP backbone of MindEye. The model performance rapidly declined, indicating that tokens processed by DFT Backbone have better mapping performance with FreMLP in the frequency domain.
@@
Table 4: Ablation experiments on modules of DFT Backbone.
@@
Table 5: Retrieval performance with different CLIP embeddings for Subject 1 on the NSD dataset.
Cerebral cortex. In order to explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s [38] method to extract the stream ROI on the NSD dataset, covering the visual cortex of nsdgenal, and trained the individual retrieval DFT Backbones for visual cortices (early, lateral, parietal, ventral). As shown in Table 6, it can be observed that the early visual cortex has the greatest impact on retrieval accuracy. Although with voxel length of 5917, Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research. They demonstrated that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving fully connected MLP Backbones are unnecessary. Interestingly, for other cortical regions, both lateral and ventral have some impacts on retrieval accuracy, while parietal shows little.
@@
Table 6: Retrieval performance with different cerebral cortex for Subject 1 on the NSD dataset.
Visualization. We visualize the CLS embeddings for LAION-5B using T-SNE in Figure 5. The figure shows that the diffusion projector successfully transformed the voxel CLS embeddings learned from contrastive learning into the image CLS embeddings, and the two kinds of embeddingss are well fused together for image-to-image retrieval on the LAION-5B dataset. We have also visualized the alignment process of voxel embeddings by contrastive learning on the NSD dataset in Appendix C.3, and learned filters of Filter Blocks in fMRI Spectrum Compression in Appendix C.5.
Figure 5: T-SNE of embeddings of 982 test fMRI by diffusion projector on LAION-5B retrieval for Subject 1. In the figure, image-clip represents image CLIP CLS embeddings, while voxel-fft and voxel-diffusion represents voxel CLS embeddings by DFT Backbone or Diffusion Projector. The diffusion projector plays a role in bringing vectors closer, enabling image-to-image retrieval on the LAION-5B dataset.
6 Limitations and Future Work
Due to inherent differences between Subjects, Lite-Mind still needs to train Subject-specific models. Although our exploratory experiments have shown that a common model can achieve comparable retrieval accuracy on four Subjects, it’s still necessary to ensure sufficient training data. Further design beyond the model level is still needed for cross-Subject models, to achieve Subject-specific few-shot or even zero-shot. In addition, Lite-Mind performs worse on Subjects with shorter voxel lengths than Subjects with longer, consistent with previous research, which lacks relevant neuroscience theoretical evidence. This provides a promising avenue for the interpretability of deep learning in neuroscience.
7 Conclusion

### v3

@@ 3.2.2. Frequency Projector
(5) $\begin{split}X^{\prime}&=({\sigma(Re(\hat{X}^{T})\mathcal{W}_{r}-Im(\hat{X}^{T})\mathcal{W}_{i}+\mathcal{B}_{r})}\\ &+j\sigma(Re(\hat{X}^{T})\mathcal{W}_{i}+Im(\hat{X}^{T})\mathcal{W}_{r}+\mathcal{B}_{i}))^{T}\end{split}$
Where $Re(\cdot)$ denotes the real part of $\hat{X}^{T}$, $Im(\cdot)$ denotes the imaginary part, $\mathcal{W}=\mathcal{W}_{r}+j\mathcal{W}_{i}$ and $\mathcal{B}=\mathcal{B}_{r}+j\mathcal{B}_{i}$. Due to the more significant features of filtered and compressed tokens in the frequency domain, we replace the final FC layer of MindEye’s MLP Backbone with FreMLP. For more theoretical proof and implementation of FreMLP, please refer to Appendix D.1. Finally, we follow IDFT $t^{\prime}\leftarrow F^{-1}(X^{\prime})$ to get final tokens $t^{\prime}$, and use voxel embedding $f$ to represent them.
3.3. Retrieval Pipeline
3.3.1. Optimization Objective
The optimization objective for the retrieval task is as follows:
@@ 3.3.1. Optimization Objective
3.3.2. Diffusion Projector
When searching on a large-scale dataset like LAION-5B, image-to-image performs better than fMRI-to-image and can find more similar images. Therefore, we used a diffusion projector to translate the disjointed voxel CLS embeddings of the DFT Backbone output, using a DALLE · 2(Ramesh et al., 2022) similar diffusion model to obtain the image CLS embeddings and retrieve it on LAION-5B: $\mathcal{V^{\prime}}=Diffusion(f)$. The Diffusion Projector consists of a transformer for image CLS embedding forecasting.
3.3.3. Contrastive Learning
@@
(7) $L_{contr}=-\frac{1}{|B|}\sum_{s=1}^{|B|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|B|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$
@@
(8) $L_{mse}=\frac{1}{|B|}\sum_{s=1}^{|B|}{||\mathcal{V}_{s}-\mathcal{V^{\prime}}_{s}||}^{2}_{2}$
@@ 3.3.4. Downstream tasks
Figure 3. Partial retrieval results of Lite-Mind on all 982 test images for Subject 1. With 12.5M DFT Backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates. The number below each image represents the similarity score. See more cases including failure retrieval in Appendix C.4.
4. Experiments
4.1. Dataset
Natural Scenes Dataset (NSD) is an extensive 7T fMRI dataset gathered from 8 subjects viewing images from MSCOCO-2017 dataset (Allen et al., 2022), which contains images of complex natural scenes. Participants viewed three repetitions of 10,000 images with a 7-Tesla fMRI scanner over 30–40 sessions. More details can be found on the NSD official website [^1] https://naturalscenesdataset.org. Our experiments focused on Subj01, Subj02, Subj05, and Subj07, who finished all viewing trials. Each subject’s training set comprises 8859 image stimuli and 24980 fMRI trials (with the possibility of 1-3 repetitions per image). The test set contains 982 image stimuli and 2770 fMRI trials. Responses for images with multiple fMRI trials are averaged across these trials. By applying the nsdgeneral ROI mask with a 1.8 mm resolution, we obtained ROIs for the four subjects, comprising 15724, 14278, 13039, and 12682 voxels respectively. These regions span visual areas ranging from the early visual cortex to higher visual areas. Our experimental setup is consistent with the NSD image reconstruction and retrieval articles (Takagi and Nishimoto, 2023; Ozcelik and VanRullen, 2023; Mai and Zhang, 2023; Scotti et al., 2023; Lin et al., 2022).
Generic Object Decoding (GOD) Dataset [^2] http://brainliner.jp/data/brainliner/Generic_Object_Decoding was created by Horikawa and Kamitani (Horikawa and Kamitani, 2017), consisting of fMRI recordings of five healthy subjects who were presented with images from ImageNet (Deng et al., 2009). The GOD Dataset includes 1250 distinct images selected from 200 ImageNet categories. Among these, 1200 training images are drawn from 150 categories, and 50 test images are from the remaining 50 categories. The training and test image stimuli were presented to the subjects once and 35 times respectively, resulting in 1200 and 1750 fMRI instances. We used preprocessed ROIs, encompassing voxels from the early visual cortex to higher visual areas. For each test image, the fMRI responses from different trials were averaged.
4.2. Implementation details
@@
Lin et al. (Lin et al., 2022) deep models 2.34M 11.0% 49.0%
Ozcelik… (Ozcelik and VanRullen, 2023) 257 separate linear regression models 3B 29.9% 21.4%
MindEye (Scotti et al., 2023) MLP Backbone 940M 89.6% 82.2%
MindEye (Scotti et al., 2023) MLP Backbone+Projector 996M 88.8% 84.9%
MindEye (Scotti et al., 2023) MLP Backbone+Prior 1.2B 93.4% 90.1%
@@
Lin et al. (Lin et al., 2022) - - - - 78.2% - - -
Takagi… (Takagi and Nishimoto, 2023) - - 83.0% 83.0% 76.0% 77.0% - -
Gu et al. (Gu et al., 2022) .150 .325 - - - - .862 .465
Ozcelik… (Ozcelik and VanRullen, 2023) .254 .356 94.2% 96.2% 87.2% 91.5% .775 .423
BrainCLIP (Liu et al., 2023) - - - - 86.7% 94.8% - -
MindEye (Scotti et al., 2023) .309 .323 94.7% 97.8% 93.8% 94.1 .645 .367
@@
Table 1. Main results of Lite-Mind retrieval and LAION-5B retrieval on the NSD dataset.
5. Results
5.1. fMRI/image retrieval
Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings on the test set. If a paired image embedding is retrieved, the retrieval is considered correct. fMRI retrieval is the opposite process mentioned above. Note that there are many semantically and visually similar images on the NSD test set, considered to be similar in CLIP space. Whether the model can correctly retrieve as MindEye tests the fine-grained alignment ability to the image. For test retrieval, we adhered to the identical methodology as Lin et al. (Lin et al., 2022) and MindEye (Scotti et al., 2023) to compute the retrieval metrics presented in Table 1(a). For each test sample, we randomly selected 299 images from the remaining 981 images in the test set and calculated the cosine similarity between the voxel embeddings and 300 images. The retrieval accuracy refers to the proportion of successful retrieval of corresponding images in the 982 voxel embeddings of the test set. We adjusted the random number seed of 30 randomly selected images to average the accuracy of all samples. The experimental results are shown in Table 1(a). We also conducted retrieval experiments on the remaining three subjects to demonstrate the universality of DFT Backbone. The detailed results and discussions are in Appendix B.2.
5.2. LAION-5B retrieval
@@
Although the retrieval scale on NSD is large enough, we can still expand it to larger datasets, such as LAION-5B (Schuhmann et al., 2022). The final layer CLIP ViT-L/14 CLS embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client (Beaumont, 2022). For LAION-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS embeddings of CLIP. For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS embeddings, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP). Unlike MindEye, in the LAION-5B retrieval process, we used the Diffusion Projector to conduct image-to-image retrieval.
@@
To compare the performance of large-scale dataset retrieval, we use metrics of image reconstruction for evaluation with previous image reconstruction methods. The experimental results are shown in Table 1(b). The specific evaluation metrics in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstruction images; SSIM is structural similarity index metric (Wang et al., 2004); EfficientNetB1(”Eff”) (Tan and Quoc, [n. d.]) and SwAV-ResNet50(”SwAV”) (Caron et al., 2020) refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
@@
Figure 4. Retrieval results of Lite-Mind from LAION-5B for Subject 1. The left column marked by a red box in every two columns represents the original image seen by the Subject, and the right column represents the image retrieved from LAION-5B.
@@
CADA-VAE (Schonfeld et al., 2019) V&T - 6.31 35.70 6.45 40.12 17.74 54.34 12.17 36.64 7.45 35.04
MVAE (Wu and Goodman, 2018) V&T - 5.77 31.51 5.40 38.46 17.11 52.46 14.02 40.90 7.89 34.63
MMVAE (Shi et al., 2019) V&T - 6.63 38.74 6.60 41.03 22.11 56.28 14.54 42.45 8.53 38.14
MoPoE-VAE (Loshchilov and Hutter, 2017) V&T - 8.54 44.05 8.34 48.11 22.68 61.82 14.57 58.51 10.45 46.40
BraVL (Du et al., 2023) V&T - 9.11 46.80 8.91 48.86 24.00 62.06 15.08 60.00 12.86 47.94
BrainClip-VAE (Liu et al., 2023) V&T Text 8.00 42.00 24.00 52.00 20.00 58.00 20.00 58.00 20.00 46.00
BrainClip-VAE (Liu et al., 2023) V&T CoOp 14.00 54.00 20.00 60.00 21.33 64.67 16.67 66.67 18.00 52.00
MindEye-211M (Scotti et al., 2023) V Text 6.00 40.00 20.00 60.00 28.00 70.00 14.00 54.00 10.00 64.00
@@
Table 2. Zero-shot visual stimulus classification on the GOD dataset. The test set contains 50 categories that have no overlapping with the training set (top-1 chance=2.0%). The results for CADA-VAE, MVAE, MMVAE, MoPoE-VAE, and BraVL are taken from (Du et al., 2023). MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B.3 for retrieval performance on the GOD dataset).
Experiments show that our DFT Backbone performs better than MindEye in high-level metrics, as shown in Table 1(b), although the backbone only has 8M parameters (MindEye-CLS backbone still has 125M. Note that both MindEye and Lite-Mind perform LAION-5B retrieval based on CLS alignment, resulting in a smaller number of parameters compared to full models). The results indicate that an efficient DFT backbone can also assist downstream retrieval tasks. The retrieval visualization results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval performance and visualization results of each Subject are shown in Appendix B.2.
5.3. GOD zero-shot classification
Since fMRI data are much fewer and shorter than those of the NSD dataset usually, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify Lite-Mind’s generalization ability. The voxel lengths collected by Subjects on the GOD dataset range from 4133 to 4643, approximately 30% of NSD, and there are only 1200 fMRI-image pairs on the training dataset (4.8% of NSD).
We conduct zero-shot classification by using fMRI-to-image retrieval to find the specific image and using simple prompt text templates ”An image of [class]” to obtain the category of retrieved images(CLIP has a classification accuracy of 76.2% on ImageNet without pre-training on it, so it can be seen as zero-shot classification). We train our Lite-Mind-15.5M to align voxels to images on the GOD dataset. We also replicated the MLP backbone and projector of MindEye, changing the residual block of $4096\times 4096$ to $1024\times 1024$ to accommodate the reduction of voxel length(We empirically found that the original $4096\times 4096$ performed even worse), and MindEye still has 210M parameters. As shown in Table 2, the experimental results indicate that our proposed Lite-Mind still guarantees robustness and establishes a new state-of-the-art for zero-shot classification on the GOD dataset without complex text prompt templates or auxiliary text representation training. On the contrary, due to the sudden decrease in dataset size and voxel length, the performance of MindEye significantly decreases, indicating excessive reliance on training data size for voxel value-wise mappings with large MLP Backbone. More retrieval and classification results are provided in Appendix B.3.
5.4. Ablations and visualization
@@
Architectural Analysis. We trained multiple DFT Backbone models with different depths to evaluate the impact of the number of layers of fMRI Spectrum Compression on retrieval efficiency(Table 3). As the Filter Block layers gradually deepen, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual. We interestingly found that the accuracy of brain retrieval tends to converge faster than the image retrieval and only 6 layers of Filter Blocks are needed to approach the highest value. We have also conducted more ablation experiments on hyper-parameters to demonstrate the robustness of Lite-Mind. Refer to Appendix B.4 for it.
@@
Table 3. Ablation experiments on fMRI Spectrum Compression revealed a positive correlation between the depth of the Filter Blocks and enhanced forward retrieval accuracy.
Module Analysis. We conducted ablation experiments on two main modules of DFT Backbone in Table 4. We completely removed the Filter blocks and aligned tokens directly to the image embeddings with frequency projector, achieving a forward retrieval accuracy of 65.8%, indicating the importance of frequency domain filtering for denoising and compression. Then we replaced FreMLP with the FC layer in the real domain, where tokens are flattened into a tensor for alignment, similar to the last layer of the MLP backbone of MindEye. The model performance rapidly declined, indicating that tokens processed by DFT Backbone have better mapping performance with FreMLP in the frequency domain.
@@
Table 4. Ablation experiments on modules of DFT Backbone.
@@
Table 5. Retrieval performance with different CLIP embeddings for Subject 1 on the NSD dataset.
Cerebral cortex. To explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s (Takagi and Nishimoto, 2023) method to extract the stream ROI on the NSD dataset, covering the visual cortex of nsdgenal, and trained the individual retrieval DFT Backbones for visual cortices (early, lateral, parietal, ventral). As shown in Table 6, it can be observed that the early visual cortex has the greatest impact on retrieval accuracy. Although with a voxel length of 5917, Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research. They demonstrated that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving that fully connected MLP Backbones are unnecessary. Interestingly, for other cortical regions, both lateral and ventral have some impacts on retrieval accuracy, while parietal shows little.
Visualization. We visualize the CLS embeddings for LAION-5B using T-SNE in Figure 5. The figure shows that the diffusion projector successfully transformed the voxel CLS embeddings learned from contrastive learning into the image CLS embeddings, and the two kinds of embeddings are well fused for image-to-image retrieval on the LAION-5B dataset. We have also visualized the alignment process of voxel embeddings by contrastive learning on the NSD dataset in Appendix C.3 and learned filters of Filter Blocks in fMRI Spectrum Compression in Appendix C.5.
Figure 5. T-SNE of embeddings of 982 test fMRI by diffusion projector on LAION-5B retrieval for Subject 1. In the figure, image-clip represents image-clip CLS embeddings, while voxel-fft and voxel-diffusion represent voxel CLS embeddings by DFT Backbone or Diffusion Projector. The diffusion projector plays a role in bringing vectors closer, enabling image-to-image retrieval on the LAION-5B dataset.
@@
Table 6. Retrieval performance with different cerebral cortex for Subject 1 on the NSD dataset.
6. Limitations and Future Work
Due to inherent differences between Subjects, Lite-Mind still needs to train Subject-specific models. Although our exploratory experiments have shown that a common model can achieve comparable retrieval accuracy on four Subjects, it’s still necessary to ensure sufficient training data. Further design beyond the model level is still needed for cross-Subject models, to achieve Subject-specific few-shot or even zero-shot. In addition, Lite-Mind performs worse on Subjects with shorter voxel lengths than on Subjects with longer, consistent with previous research, which lacks relevant neuroscience theoretical evidence. This provides a promising avenue for the interpretability of deep learning in neuroscience.
7. Conclusion